# Massive tree-level splitting functions beyond kinematical limits

### v1

@@ VI Conclusions
We presented a re-computation of the massive QCD one-to-three parton splitting functions and a significantly more compact form of the results compared to existing literature. This will reduce evaluation time and increase numerical stability in practical applications. Using a recently introduced method for the decomposition of splitting functions into scalar dipole radiator functions, lower-order results, and pure splitting remainders, we derived a representation that makes the factorization into leading-order results transparent. The pure splitting remainders exhibit at most sub-leading singularities in the (double-)soft gluon limits and in the quasi-collinear limits. We expect our novel technique to provide a path towards higher-order computations which do not require sectorization or other forms of overlap removal for different singular phase-space configurations. At the same time, the splitting functions derived here will be helpful for higher logarithmic resummation of soft gluon effects on QCD observables. These methods will help drive improvements in the modeling of heavy quark jet formation, leading to better precision in analyses of top quark and electroweak boson production at the LHC.

### v2

@@ VI Conclusions
Using a recently introduced method for the decomposition of splitting functions into scalar dipole radiators, lower-order expressions, and pure splitting remainders, we derived a representation of the massive one-to-three-parton splittings that makes their factorization manifest. The massive two-gluon radiators used in this context were computed in a process-independent form for the first time. The pure splitting remainders exhibit at most sub-leading singularities in the (double-)soft gluon limits and in the quasi-collinear limits. Kinematical limits are, however, not used in any of our calculations, which implies that the results remain valid across the complete physical phase space.
As a byproduct of our analysis, we derived the massive QCD one-to-three parton splitting functions in a significantly more compact form than existing literature. This will reduce evaluation time and increase numerical stability in practical applications. We have also obtained the first independent cross-check of the massive, triple-collinear splitting functions that do not exist in the SCET-based approach of Craft et al. (2024), in particular Eqs. (5.9) and (5.10) of Dhani et al. (2023).
We expect the novel techniques introduced in this manuscript to provide a path towards higher-order computations which do not require sectorization or other forms of overlap removal for different singular phase-space configurations. At the same time, the splitting functions derived here will be helpful for higher logarithmic resummation of soft gluon effects on QCD observables. These methods will help drive improvements in the modeling of heavy quark jet formation, leading to better precision in analyses of top quark and electroweak boson production at the LHC.